# LSM Trees in Adversarial Environments

### v1

@@ Abstract.
The Log Structured Merge (LSM) Tree is a popular choice for key-value stores that focus on optimized write throughput while maintaining performant, production-ready read latencies. To optimize read performance, LSM stores rely on a probabilistic data structure called the Bloom Filter (BF). In this paper, we focus on adversarial workloads that lead to a sharp degradation in read performance by impacting the accuracy of BFs used within the LSM store. Our evaluation shows up to $800\%$ increase in the read latency of lookups for popular LSM stores. We define adversarial models and security definitions for LSM stores. We implement adversary resilience into two popular LSM stores, LevelDB and RocksDB. We use our implementations to demonstrate how performance degradation under adversarial workloads can be mitigated.
LSM stores, Storage Security, Bloom Filters, Cryptography
copyright: acmcopyrightconference: ; ; journalyear: 2025ccs: Information systems Key-value storesccs: Security and privacy Database and storage securityccs: Theory of computation Theory of database privacy and security
@@ 1. Introduction
LSM Stores. A large number of modern key-value stores are based on Log-structured Merge Trees (LSM Trees) (Kaushik and Sarkar, 2024; Dayan et al., 2017). LSM trees allow for write-optimized (Dayan et al., 2017), highly configurable (Thakkar et al., 2024) storage while being relatively simple to implement. These qualities have made them common as the storage engine for a large number of commercial and open-source database systems including Google’s LevelDB (Google, 2025), Meta’s RocksDB (Meta, 2025), MongoDB’s WiredTiger (MongoDB, 2025) and others (Kaushik and Sarkar, 2024; Dayan et al., 2017; Sears and Ramakrishnan, 2012; Apache, 2025b, a).
Adversarial Performance. LSM stores use Bloom Filters (Bloom, 1970; Broder and Mitzenmacher, 2003) to reduce unnecessary disk access. This strategy depends on Bloom Filters maintaining a low false positive rate (FPR). In normal workloads, this works well as Bloom Filters filter non-existent keys, so the LSM store does not have to look on disk (Dayan et al., 2017). An adversary can strategically insert keys into an LSM store that saturate its Bloom Filters, drastically raising their FPR. In this case, lookups for non-existing keys (called zero-result lookups (Dayan et al., 2017)) require multiple disk accesses, increasing query latency by up to 800% according to our experiments. An adversary only needs a modest number of well-chosen insertions to render Bloom Filters in LSM stores ineffective. This motivates the need for adversarially resilient mechanisms in LSM store.
Contributions. Our contributions are as follows.
@@
Related Work. There have been many recent efforts to rigorously define a game-based (Naor and Eylon, 2019; Naor and Oved, 2022; Bishop and Tirmazi, 2024; Clayton et al., 2019) and simulator-based (Filić et al., 2022, 2024; Tirmazi, 2025) adversarial model for probabilistic data structures such as the Bloom Filter and the Learned Bloom Filter. Our work builds upon these insights by applying them to LSM stores and designing countermeasures tailored to real-world storage systems. Prior work on LSM store optimizations including (Dayan et al., 2017; Kaushik and Sarkar, 2024; Thakkar et al., 2024; Dayan et al., 2022; Huynh et al., 2022) has primarily focused on improving performance under non-adversarial workloads. To the best of our knowledge, this is the first work to rigorously define an adversarial model for LSM stores and propose concrete, provably secure LSM store constructions.
@@ 2. Preliminaries
LSM stores. An LSM Tree consists of $\mathcal{N}$ levels (Dayan et al., 2017). $L_{i}$ indicates level $i$ in the LSM Tree. $L_{0}$ is typically an in-memory buffer, while the remaining levels $[L_{1},\cdots,L_{\mathcal{N}}]$ consist of data in secondary storage including SSDs, Hard Disks, and external storage nodes connected over the network. We formalize the general syntax and behavior of an LSM store. Given an LSM store, $\Lambda$, we denote the set of public parameters of an LSM store by $\Phi$. The public parameters contain all the knobs relevant to the LSM store implementation including the number of levels in the LSM Tree as well as the knobs of any auxiliary data structures (discussed below) used by the LSM store. We denote the elements stored in $\Lambda$ by $\mathcal{L}_{\Lambda}$ (a list). We denote the state of $\Lambda$ by $\sigma\in\Sigma$ where $\Sigma$ is the space of all possible states of $\Lambda$. $\Lambda$ can store elements from any finite domain $\mathfrak{D}$, where $\mathfrak{D}=\cup_{l=0}^{L}\{0,1\}^{l}$ for any natural number $L\in\mathbf{N}$. An LSM store, $\Lambda$ consists of three algorithms.
@@
LSM Axioms. We define two axioms to reason about the performance of an LSM store, the Axiom of Cost (Axiom 1) and the Axiom of Recency (Axiom 2). Our axioms hold for many popular LSM store implementations including LevelDB (Google, 2025), RocksDB (Meta, 2025), Monkey (Dayan et al., 2017), WiredTiger (MongoDB, 2025), HBase (Apache, 2025b) and Cassandra (Apache, 2025a).
@@
Using these axioms we explain store design and performance.
Zero-result Lookups. In an LSM Tree with $L_{\mathcal{N}}$ levels, when any key $k$ is queried, the LSM store begins at the in-memory buffer at $L_{0}$ and traverses from $L_{0}$ to $L_{\mathcal{N}}$ in ascending order. If the LSM store finds a matching key at level $L_{i}$, the query returns early (Dayan et al., 2017). Entries at higher levels are superseded by more recent entries at lower levels thanks to Axiom 2, so looking further is unnecessary. For a key whose most recent entry is present in run $R_{i,x}$, the number of runs an LSM store needs to probe is at most $\sum_{l=0}^{i}\sum_{r=0}^{N_{R}(L_{l})}E_{C}(R_{l,r})$ where $E_{C}(r)$ is the expected I/O cost of probing run $r$.
@@
Auxiliary Data Structures. For each run $R_{i,x}$ in an LSM Tree, modern LSM stores store two auxiliary data structures in main memory (Dayan et al., 2017): a Bloom Filter (BF) (Bloom, 1970) and an array of fence pointers (Huynh et al., 2022; Dayan et al., 2017). The fence pointers contain min/max information for the disk pages that store run $R_{i,x}$ (Meta, 2025; Google, 2025). In a point query, the LSM store uses this array of fence pointers when traversing a level to figure out which disk page to read when searching for a key (Dayan et al., 2017). The LSM store only has to read one disk page for each run.
@@
3. Adversarial Workloads
In this section, we discuss the performance degradation of LSM stores under adversarial workloads.
@@ 3. Adversarial Workloads
Performance Degradation. The key idea for our adversarial workloads is to target the false positive rate (FPR) of the per-run Bloom Filters implemented in an LSM store. Since most popular LSM stores (Google, 2025; Meta, 2025; Dayan et al., 2017) use BFs with non-cryptographic hashes, such attacks are feasible and inexpensive. Even if an adversary is allowed a very small bound on insertions, they can still fully insert enough elements to saturate the BFs. For a Standard Bloom Filter implementation with a memory budget of $m_{\dagger}$ bits and $k_{\dagger}$ hashes, only $\frac{m_{\dagger}}{k_{\dagger}}$ adversarial insertions are required to make the BF’s FPR become $1$. Figure 1 shows the impact of Bloom Filter FPR on the zero-result lookup latency for LevelDB (Google, 2025) and RocksDB (Meta, 2025). We insert $10$M keys in batch sizes of $10$K. We benchmark $50$K uniformly randomly sampled zero-result look-ups for each experiment. Between the insertion stage and the lookup stage, we save and re-open the store to mitigate the effects of caching. As BF FPR increases, the lookup latency of LevelDB and RocksDB becomes $2$x and $8$x respectively.
Feasibility of Attacks. We first give an overview of how the standard implementation of a Bloom Filter first suggested in (Bloom, 1970) and used in LevelDB (Google, 2025) and RocksDB (Meta, 2025) works. A Standard Bloom Filter (SBF) construction is a zero-initialized array of $m_{\dagger}$ bits (Broder and Mitzenmacher, 2003; Gerbet et al., 2015) and requires a family of $k_{\dagger}$ independent hash functions, $\,\,h_{i,m}:\mathfrak{D}\mapsto[m_{\dagger}]$ for all $i\in[k_{\dagger}]$(Broder and Mitzenmacher, 2003; Tirmazi, 2025). Upon setup, For each element $x\in S$ ($S$ is the set being encoded by the Bloom Filter), the bits $h_{i}(x)$ are set to $1$ for $i\in[k_{\dagger}]$. When querying an element $x$, we return true if all $h_{i}(x)$ map to bits that are set to $1$. If there exists an $h_{i}(x)$ that maps to a bit that is 00, we return false.
@@
4. Adversarial Model
We define a self-contained game-based (Naor and Eylon, 2019; Naor and Oved, 2022; Bishop and Tirmazi, 2024) adversarial model for LSM stores that is restricted to adversaries that target an LSM store’s Bloom Filters. For cryptography-focused readers, we also define a more general simulator-based (Lindell, 2016) adversarial model for any class of adversaries in Section A.
Adversary Games. We will be assuming our adversary is efficient. We will assume that the adversary works in non-uniform probabilistic polynomial time (n.u. p.p.t). This standard notion is used to model efficient adversaries in cryptography literature (Pass and Shelat, 2010). Moreover, we will restrict our model to $\mathbb{A}_{\text{BF}}$, the set of efficient adversaries that target the Bloom Filters in the LSM store. We discuss other adversarial targets in Appendix C. To capture the notion of adversary resilience in an LSM store setting, we propose a game inspired by prior work on game-based adversary models for other probabilistic data structures (Naor and Eylon, 2019; Naor and Oved, 2022). We first define the notion of a Bloom Filter in our adversarial game.
@@ 4. Adversarial Model
Security Definition. We now define our notion of the security of an LSM store $\Lambda$ against Bloom Filter targeting adversaries $\mathbb{A}_{\text{BF}}$ using the Smash-LSM game.
@@ 5. Adversary Resilience
We construct an adversary-resilient implementation of LevelDB (Google, 2025) and RocksDB (Meta, 2025). Our secure construction is simple and easily pluggable in any popular LSM store. Instead of writing and reading keys directly, we simply patch the LSM store API to maintain a keyed pseudorandom permutation (PRP) and read/write the permuted value of the key. The values are kept unchanged. This does not affect the correctness of the LSM store. As an example, assume we use a PRP that permutes $a$ to $x$ and $b$ to $y$. Inserting two key-value pairs $(a,v_{1})$ and $(b,v_{2})$ will instead insert $(x,v_{1})$ and $(y,v_{2})$. When key $a$ is queries, since our PRP is consistent, it is mapped again to $x$ and $\Lambda$ returns the correct corresponding value $v_{1}$. PRPs are bijections therefore the keys themselves can also be recovered by running the inverse permutation. We show in Appendix B that our construction is provably secure against Bloom Filter targeting adversaries. We discuss our main result here.
@@
Appendix B (Thm B.6) contains a rigorous proof. As a proof sketch, applying a keyed PRP to LSM store keys before insertion and query ensures that adversaries cannot control key placements to force high false positive rates. Correctness is due to PRP bijectivity: queries map consistently to transformed keys, retrieving the correct values. Security follows from PRP indistinguishability. If an adversary could significantly increase false positives, they could distinguish the PRP from random, which contradicts the existence of PRPs. We need extra $\lambda$ bits of memory to store the PRP’s secret key.
Implementation & Experimental Setup. We use the experimental setup discussed here for all experiments in this work. We use an Apple M2 processor with $8$ cores, $8$ GB of Memory, a $128$ KB L1 cache, a $1$ MB L2 cache, and $256$ GB of SSD storage out of which $245$ GB is usable storage. We evaluated the most recent versions of RocksDB ($9.10$) and LevelDB ($1.23$). We run each experiment $5$ times and display the median, the error bars indicate standard deviations. We implement security as an easily pluggable module written in C++ 17 for both LevelDB and RocksDB. Our implementation relies on the hardness of AES. Our implementation uses the AES-128 implementation provided by OpenSSL v3.4.0 in CTR mode. We use a $16$ byte secret key for our experiments. We have released our implementation and evaluation as open-source software [^1] The codebase of this work is available at https://github.com/jadidbourbaki/db_attacks.
Performance. We compare our adversary-resilient implementations to untouched releases of LevelDB and RocksDB under adversarial workloads. Figures4 shows performance under the same workload as Section 3. For zero-result lookups, our implementations reduce latency by 60% and 78% for LevelDB and RocksDB respectively. Our implementations are slower for workloads with existing keys (where not all the runs need to be probed, in expectation) by 30% and 6% for LevelDB and RocksDB respectively. We also measured the latency for $50$ K uniformly random inserts with our implementation and saw a median latency increase of 39% and 49% for LevelDB and RocksDB respectively. These increases are due to the extra compute overhead of AES for every key. This is a constant compute cost that can potentially be reduced by using a hardware-implemented version of AES such as AES-NI (Naor and Eylon, 2019) for Intel processors. This overhead also is not a major concern when I/O dominates lookup performance, which is frequently the case for LSM stores (Dayan et al., 2017).
@@
Figure 4. Zero-result lookup performance on a uniformly random query workload for LevelDB (top left) and RocksDB (top right) with adversarial resilience and existing query lookup performance on the same workload for LevelDB (bottom left) and RocksDB (bottom right) with adversarial resilience.

### v2

@@ Abstract.
The Log Structured Merge (LSM) Tree is a popular choice for key-value stores focusing on optimized write throughput while maintaining performant, production-ready read latencies. LSM stores rely on a probabilistic data structure called the Bloom Filter (BF) to optimize read performance. In this paper, we focus on adversarial workloads that lead to a sharp degradation in read performance by impacting the accuracy of BFs used within the LSM store. Our evaluation shows up to $800\%$ increase in the read latency of lookups for popular LSM stores. We define adversarial models and security definitions for LSM stores. We implement adversary resilience into two popular LSM stores, LevelDB and RocksDB. We use our implementations to demonstrate how performance degradation under adversarial workloads can be mitigated.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/jadidbourbaki/db_attacks
@@ 1. Introduction
A large number of modern key-value stores are based on Log-structured Merge Trees (LSM Trees) (Kaushik and Sarkar, 2024; Dayan et al., 2017). LSM trees allow for write-optimized (Dayan et al., 2017), highly configurable (Thakkar et al., 2024) storage while being relatively simple to implement. These qualities have made them common as the storage engine for a large number of commercial and open-source database systems including LevelDB (Google, 2025) from Google, bLSM (Sears and Ramakrishnan, 2012) and cLSM (Golan-Gueta et al., 2015) from Yahoo, RocksDB (Meta, 2025) and Cassandra (Apache, 2025a) from Meta, WiredTiger (MongoDB, 2025) from MongoDB, Monkey (Dayan et al., 2017) from Harvard’s Data Systems Lab, Apache HBase (Apache, 2025b),  (DGraph, 2025) from DGraph and many others (Kaushik and Sarkar, 2024; Dayan and Idreos, 2019).
LSM stores use Bloom Filters (Bloom, 1970; Broder and Mitzenmacher, 2003) to reduce unnecessary disk access. This strategy depends on Bloom Filters maintaining a low false positive rate (FPR). In normal workloads, this works well as Bloom Filters filter non-existent keys, so the LSM store does not have to look on disk (Dayan et al., 2017). An adversary can strategically insert keys into an LSM store that saturate its Bloom Filters, drastically raising their FPR. In this case, lookups for non-existing keys (called zero-result lookups (Dayan et al., 2017)) require multiple disk accesses, increasing query latency by up to 800% according to our experiments. An adversary only needs a modest number of well-chosen insertions to render Bloom Filters in LSM stores ineffective. This motivates the need for adversarially resilient mechanisms in LSM store.
We make the following contributions in this paper.
@@ 2. Preliminaries
In this section, we first go over notation and rigorously define an LSM store. We then define a set of axioms that hold for a large number of popular LSM store implementations. Lastly, we use this axiomatic framework to introduce multiple aspects of LSM store design and performance include the performance of zero-result lookups and the use of auxiliary data structures.
Notation.
@@
2.1. LSM stores
An LSM Tree consists of $\mathcal{N}$ levels (Dayan et al., 2017). $L_{i}$ indicates level $i$ in the LSM Tree. $L_{0}$ is typically an in-memory buffer, while the remaining levels $[L_{1},\cdots,L_{\mathcal{N}}]$ consist of data in secondary storage including SSDs, Hard Disks, and external storage nodes connected over the network. We formalize the general syntax and behavior of an LSM store. Given an LSM store, $\Lambda$, we denote the set of public parameters of an LSM store by $\Phi$. The public parameters contain all the knobs relevant to the LSM store implementation including the number of levels in the LSM Tree as well as the knobs of any auxiliary data structures (discussed below) used by the LSM store. We denote the elements stored in $\Lambda$ by $\mathcal{L}_{\Lambda}$ (a list). We denote the state of $\Lambda$ by $\sigma\in\Sigma$ where $\Sigma$ is the space of all possible states of $\Lambda$. $\Lambda$ can store elements from any finite domain $\mathfrak{D}$, where $\mathfrak{D}=\cup_{l=0}^{L}\{0,1\}^{l}$ for any natural number $L\in\mathbf{N}$. An LSM store, $\Lambda$ consists of three algorithms.
@@ 2.1. LSM stores
Symbol Description
General Notation
$\mathfrak{D}$ Finite domain of elements
$S$ A set of elements
$x\mathrel{{\leftarrow}\vbox{\hbox{\tiny\$}}}S$ $x$ sampled randomly from $S$
$|S|$ Cardinality of set $S$
$[n]$ Set $\{1,\dots,n\}$
LSM Store Notation
$\sigma\in\Sigma$ State of LSM store
$\Lambda$ An LSM store
$\Phi$ Public parameters
$L_{i}$ Level $i$ in LSM Tree
$R_{i,x}$ Run $x$ in level $L_{i}$
$N_{R}(L_{i})$ Number of runs in $L_{i}$
$E_{C}(L_{i})$ Expected access cost in $L_{i}$
$(k,v)$ Key-value pair
$C_{r}(\Phi)$ Construction algorithm
$Q(k,\sigma)$ Query operation for key $k$
$I_{r}((k,v),\sigma)$ Insert operation
$\perp$ Tombstone (deletion marker)
Bloom Filter Notation
$\Pi=({C_{\dagger}}_{r},Q_{\dagger})$ Bloom Filter (BF) construction/query
$\sigma_{\dagger}$ Internal state of BF
$m_{\dagger}$ BF bit array size
$k_{\dagger}$ Number of BF hash functions
$h_{i}(k)$ $i^{th}$ BF hash function
$P_{B}(R_{i,x})$ BF false positive probability
Adversarial Model Notation
$\mathcal{A}\in\mathbb{A}$ Computationally-bound adversary
$\mathcal{A}\in\mathbb{A}_{BF}$ BF-targeting adversary
$\lambda$ Security parameter
$t$ Number of adversary queries
Smash-Lsm Adversarial game for LSM stores
Smash-Bloom Adversarial game for Bloom Filters
$F_{\kappa}$ Keyed pseudorandom permutation
$\epsilon$ False positive/security bound
Oracle Notation
${\mathscr{O}_{Q}(k)}$ returns $\top$ if $\exists$ a BF $\Pi$ in LSM $\Lambda$ s.t $\Pi(k)=\top$
$\mathscr{O}_{R}$ Returns state $\sigma_{\dagger}$ of each BF $\Pi_{i}$ in LSM $\Lambda$
$\mathscr{O}_{I}(k,v)$ Inserts $(k,v)$ into LSM $\Lambda$
$\mathscr{O}_{C}(\Phi)$ Constructs LSM $\Lambda$ with parameters $\Phi$
${\mathscr{O}_{\dagger}}_{Q}(k)$ Queries BF $\Pi$ for $k$
${\mathscr{O}_{\dagger}}_{R}$ Returns state $\sigma_{\dagger}$ of BF $\Pi$
Table 1. Mathematical notation used in the paper
2.2. LSM Axioms
We define two axioms to reason about the performance of an LSM store, the Axiom of Cost (Axiom 1) and the Axiom of Recency (Axiom 2). Our axioms hold for many popular LSM store implementations including LevelDB (Google, 2025), RocksDB (Meta, 2025), Monkey (Dayan et al., 2017), WiredTiger (MongoDB, 2025), HBase (Apache, 2025b) and Cassandra (Apache, 2025a).
@@ 2.2. LSM Axioms
Using these axioms, we can explain LSM store design and performance.
2.3. Zero-result Lookups.
In an LSM Tree with $L_{\mathcal{N}}$ levels, when any key $k$ is queried, the LSM store begins at the in-memory buffer at $L_{0}$ and traverses from $L_{0}$ to $L_{\mathcal{N}}$ in ascending order. If the LSM store finds a matching key at level $L_{i}$, the query returns early (Dayan et al., 2017). Entries at higher levels are superseded by more recent entries at lower levels thanks to Axiom 2, so looking further is unnecessary. For a key whose most recent entry is present in run $R_{i,x}$, the number of runs an LSM store needs to probe is at most $\sum_{l=0}^{i}\sum_{r=0}^{N_{R}(L_{l})}E_{C}(R_{l,r})$ where $E_{C}(r)$ is the expected I/O cost of probing run $r$.
@@
2.4. Auxiliary Data Structures
For each run $R_{i,x}$ in an LSM Tree, modern LSM stores store two auxiliary data structures in main memory (Dayan et al., 2017): a Bloom Filter (BF) (Bloom, 1970) and an array of fence pointers (Huynh et al., 2022; Dayan et al., 2017). The fence pointers contain min/max information for the disk pages that store run $R_{i,x}$ (Meta, 2025; Google, 2025). In a point query, the LSM store uses this array of fence pointers when traversing a level to figure out which disk page to read when searching for a key (Dayan et al., 2017). The LSM store only has to read one disk page for each run.
@@ 2.4. Auxiliary Data Structures
3. Adversarial Model
In this section, we first discuss the performance degradation caused by adversarial attacks and the feasibility of such attacks. We then introduce a game-based adversarial model for LSM stores and propose a security definition for LSM stores.
@@ 3. Adversarial Model
3.1. Performance Degradation
The key idea for our adversarial workloads is to target the false positive rate (FPR) of the per-run Bloom Filters implemented in an LSM store. Since most popular LSM stores (Google, 2025; Meta, 2025; Dayan et al., 2017) use BFs with non-cryptographic hashes, such attacks are feasible and inexpensive. Even if an adversary is allowed a very small bound on insertions, they can still fully insert enough elements to saturate the BFs. For a Standard Bloom Filter implementation with a memory budget of $m_{\dagger}$ bits and $k_{\dagger}$ hashes, only $\frac{m_{\dagger}}{k_{\dagger}}$ adversarial insertions are required to make the BF’s FPR become $1$. Figure 1 shows the impact of Bloom Filter FPR on the zero-result lookup latency for LevelDB (Google, 2025) and RocksDB (Meta, 2025). We insert $10$M keys in batch sizes of $10$K. We benchmark $50$K uniformly randomly sampled zero-result look-ups for each experiment. Between the insertion stage and the lookup stage, we save and re-open the store to mitigate the effects of caching. As BF FPR increases, the lookup latency of LevelDB and RocksDB becomes $2$x and $8$x respectively.
3.2. Feasibility of Attacks
We first give an overview of how the standard implementation of a Bloom Filter first suggested in (Bloom, 1970) and used in LevelDB (Google, 2025) and RocksDB (Meta, 2025) works. A Standard Bloom Filter (SBF) construction is a zero-initialized array of $m_{\dagger}$ bits (Broder and Mitzenmacher, 2003; Gerbet et al., 2015) and requires a family of $k_{\dagger}$ independent hash functions, $\,\,h_{i,m}:\mathfrak{D}\mapsto[m_{\dagger}]$ for all $i\in[k_{\dagger}]$(Broder and Mitzenmacher, 2003; Tirmazi, 2025). Upon setup, For each element $x\in S$ ($S$ is the set being encoded by the Bloom Filter), the bits $h_{i}(x)$ are set to $1$ for $i\in[k_{\dagger}]$. When querying an element $x$, we return true if all $h_{i}(x)$ map to bits that are set to $1$. If there exists an $h_{i}(x)$ that maps to a bit that is 00, we return false.
@@ 3.2. Feasibility of Attacks
[…]
Three graphs showing how a Bloom Filter is initialized, performs inserts, and performs queries.
@@
[…]
A line graph showing a brute force algorithm taking O(seconds) to saturate Bloom Filters
Figure 4. Zero-result lookup performance on a uniformly random query workload for LevelDB ( left) and RocksDB (right) with adversarial resilience.
3.3. Game-based Model
We define a self-contained game-based (Naor and Eylon, 2019; Naor and Oved, 2022; Bishop and Tirmazi, 2024) adversarial model for LSM stores that is restricted to adversaries that target an LSM store’s Bloom Filters. For cryptography-focused readers, we also define a more general simulator-based (Lindell, 2016) adversarial model for any class of adversaries in Appendix A. We will be assuming our adversary is efficient. We will assume that the adversary works in non-uniform probabilistic polynomial time (n.u. p.p.t). This standard notion is used to model efficient adversaries in cryptography literature (Pass and Shelat, 2010). Moreover, we will restrict our model to $\mathbb{A}_{\text{BF}}$, the set of efficient adversaries that target the Bloom Filters in the LSM store. We discuss other adversarial targets in Appendix LABEL:sec:adversarial_targets. To capture the notion of adversary resilience in an LSM store setting, we propose a game inspired by prior work on game-based adversary models for other probabilistic data structures (Naor and Eylon, 2019; Naor and Oved, 2022). We first define the notion of a Bloom Filter in our adversarial game.
@@ 3.3. Game-based Model
3.4. Security Definition
We now define our notion of the security of an LSM store $\Lambda$ against Bloom Filter targeting adversaries $\mathbb{A}_{\text{BF}}$ using the Smash-LSM game.
@@ 4. Adversary Resilience
We construct an adversary-resilient implementation of LevelDB (Google, 2025) and RocksDB (Meta, 2025). Our secure construction is simple and easily pluggable in any popular LSM store. Instead of writing and reading keys directly, we simply patch the LSM store API to maintain a keyed pseudorandom permutation (PRP) and read/write the permuted value of the key. The values are kept unchanged. This does not affect the correctness of the LSM store. As an example, assume we use a PRP that permutes $a$ to $x$ and $b$ to $y$. Inserting two key-value pairs $(a,v_{1})$ and $(b,v_{2})$ will instead insert $(x,v_{1})$ and $(y,v_{2})$. When key $a$ is queries, since our PRP is consistent, it is mapped again to $x$ and $\Lambda$ returns the correct corresponding value $v_{1}$. PRPs are bijections therefore the keys themselves can also be recovered by running the inverse permutation. We informally summarize our result here and then rigorously prove it below in Section 5.
@@
Proof Sketch: Applying a keyed PRP to LSM store keys before insertion and query ensures that adversaries cannot control key placements to force high false positive rates. Correctness is due to PRP bijectivity: queries map consistently to transformed keys, retrieving the correct values. Security follows from PRP indistinguishability. If an adversary could significantly increase false positives, they could distinguish the PRP from random, which contradicts the existence of PRPs. We need extra $\lambda$ bits of memory to store the PRP’s secret key.
@@
Figure 5. Existing query lookup performance on a uniformly random query workload for LevelDB (left) and RocksDB (right) with adversarial resilience.
4.1. Implementation & Experimental Setup
We use the experimental setup discussed here for all experiments in this work. We use an Apple M2 processor with $8$ cores, $8$ GB of Memory, a $128$ KB L1 cache, a $1$ MB L2 cache, and $256$ GB of SSD storage out of which $245$ GB is usable storage. We evaluated the most recent versions of RocksDB ($9.10$) and LevelDB ($1.23$). We run each experiment $5$ times and display the median, the error bars indicate standard deviations. We implement security as an easily pluggable module written in C++ 17 for both LevelDB and RocksDB. Our implementation relies on the hardness of AES. Our implementation uses the AES-128 implementation provided by OpenSSL v3.4.0 in CTR mode. We use a $16$ byte ($\lambda=128$ bits) AES secret key to parameterize the PRP. To prevent secret key leakage, production-scale deployments can rely on secret stores and ephemeral key rotation mechanisms. We have released our implementation and evaluation as open-source software.
4.2. Performance
We compare our adversary-resilient implementations to untouched releases of LevelDB and RocksDB under adversarial workloads. Figures4 and 5 show performance under the same workload as Section 3. For zero-result lookups, our implementations reduce latency by 60% and 78% for LevelDB and RocksDB respectively. Our implementations are slower for workloads with existing keys (where not all the runs need to be probed, in expectation) by 30% and 6% for LevelDB and RocksDB respectively. We also measured the latency for $50$ K uniformly random inserts with our implementation and saw a median latency increase of 39% and 49% for LevelDB and RocksDB respectively. These increases are due to the extra compute overhead of AES for every key. This is a constant compute cost that can potentially be reduced by using a hardware-implemented version of AES such as AES-NI (Naor and Eylon, 2019) for Intel processors. This overhead also is not a major concern when I/O dominates lookup performance, which is frequently the case for LSM stores (Dayan et al., 2017).
5. Security Proofs
We first define the notion of secure Bloom Filters (Def. 3.1). We use known definitions (Filić et al., 2022; Tirmazi, 2025) for Bloom Filter (BF) constructions. A BF, $\Pi$ consists of two algorithms.
Game 5.1 (Smash-Bloom).
Step 1:
Step 2:
Step 3:
Step 4:
Definition 5.0.
[…]
Definition 5.0.
Definition 5.0.
Definition 5.0.
$\leq\mathtt{negl}(n)$
Theorem 5.5.
Proof.
  1. (1)
  2. (2)
[…]
We can now prove our main result regarding the security of LSM stores.
Theorem 5.6.
Proof.
  1. (1)
    $I^{\prime}_{r}((k,v),\sigma)=C_{r}((F_{\kappa}(k),v),\sigma)$
  2. (2)
    $Q^{\prime}(x,\sigma)=Q(F_{\kappa}(x),\sigma)$.
Security Proof: We prove this by contradiction. Suppose there does exist an n.u. p.p.t. adversary $\mathcal{A}$ that can with the Smash-Lsm game with probability greater than $\varepsilon+\textrm{negl}(\lambda)$. Then by definition of oracle $\mathscr{O}_{Q}$, there exists a Bloom Filter $\Pi^{\prime}$ with the construction from Thm 5.5 for which $\mathcal{A}$ can generate false positives with probability higher than $\varepsilon+\mathrm{negl}(\lambda)$. So $\mathcal{A}$ can also win the Smash-Bloom game with a probability higher than $\varepsilon+\mathrm{negl}(\lambda)$. This is a contradiction as Thm 5.5 proves that no n.u p.p.t adversary can win the Smash-Bloom game against Bloom Filter constructions of type $\Pi^{\prime}$ with probability higher than $\varepsilon+\mathrm{negl}(\lambda)$.
6. Related Work
We discuss two areas of research relevant to our work: adversarial correctness of probabilistic data structures, and LSM store benchmarking and optimization.
Adversarial Data Structures.
There have been many recent efforts to rigorously define a game-based (Naor and Eylon, 2019; Naor and Oved, 2022; Bishop and Tirmazi, 2024; Clayton et al., 2019) and simulator-based (Filić et al., 2022, 2024; Tirmazi, 2025) adversarial model for probabilistic data structures such as the Bloom Filter and the Learned Bloom Filter. In particular, the work of Naor et. al. (Naor and Eylon, 2019; Naor and Oved, 2022) was the first to show provably secure constructions for the Bloom Filter. There is also prior work showing feasible attacks on probabilistic data structures including the Bloom Filter (Gerbet et al., 2015) and the Learned Bloom Filter (Reviriego et al., 2021). Our work builds upon these insights by applying them to LSM stores and designing countermeasures tailored to real-world storage systems.
LSM Store Benchmarking / Optimization.
KVBench (Zhu et al., 2024) is one of many works that focus on benchmarking workloads for LSM stores. Prior work on LSM store optimizations includes  (Dayan et al., 2017; Thakkar et al., 2024; Dayan et al., 2022; Huynh et al., 2022). Monkey (Dayan et al., 2017) focuses on optimizing the memory budget allocation of the Bloom Filters used in LSM stores for better query performance. Huynch et. al. (Huynh et al., 2022) use Large Language Models (LLMs) to tune the design knobs of LSM stores. All prior work discussed primarily focuses on benchmarking and improving performance under non-adversarial workloads. To the best of our knowledge, this is the first paper to rigorously define an adversarial model for LSM stores and propose concrete, provably secure LSM store constructions.